OpenStudy (anonymous):

please help my answer 11 x 4=44 8 x 5=40 44 x 40=1,760

OpenStudy (anonymous):

@ShadowLegendX

OpenStudy (anonymous):

\[4\times 8\] for the rectangle, then \(\frac{1}{2}\times 4\times 3\) for the triangle

OpenStudy (anonymous):

how do i do 1/2 x 4?

OpenStudy (anonymous):

4 x 8=32 1/2 x 4 x 3=6 32 + 6=38 is this right??? @satellite73

OpenStudy (anonymous):

@Destinymasha please check my answer